Find the sum of: 
– 50, – 200 and 300 ????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

-50, -200 and 300
(-50) + (-200) + (+300)
= (-50) + (-200) + (+200) + (+100)
= (-50) + 0 + (+100) [(-a) + (+a) = 0]
= (-50) + (+50) + (+50)
= 0 + (+50) [(-a) + (+a) = 0]
= 50

Thanks